A force f is expressed with respect to the basis e1 and e2 by equation 2(3e1+4e2) N. If the directions of  e1  makes an angle of 30 degree with f, find the vector  e2. 


1
Expert's answer
2021-01-25T13:55:22-0500

Solution

Given force magnitude


"f=6e \n1\n\u200b\t\n +8e \n2\n\u200b"

Angle between e1 and f is =30°

Using basis vector concept

"f. e_1=(6e_1+8e_2). e_1\\\\\n\u2223(6e \n1\n\u200b\t\n +8e \n2\n\u200b\t\n )\u2223|e \n1\n\u200b\t\n \u2223cos30 \n\u00b0\n =6"

Magnitude of e1 is unit.

So vector e2 can be written as

"e_2=\frac{4\sqrt{3}-6e_1}{8}=\frac{\sqrt{3}}{2}-\frac{3e_1}{4}"